18. Researchers have investigated the preferences over dinner and they have found that they take the...

18.

Researchers have investigated the preferences over dinner and they have found that they take the following form:

U = - 68 Chicken + 21 Steak + 4 Fish - 43 Costs

Costs= cost in dollars of the dinner
Chicken= amount of chicken in pounds
Steak= amount of steak in pounds
Fish= amount of fish in pounds

Given this information, how many pounds of fish is the consumer willing to give up to obtain one more pound of steak during dinner? (hint: use MRS(steak, fish) but enter your answer as a positive number)

19. If the ulitity function is: U=86x1 + 14x2

If x1=44, how much of x2 does the consumer has to consume to obtain a utility level of 6,159?
Answer #1

18. MRS(steak, fish) = -Marginal utility of steak/Marginal utility of fish = (entering positive number only)

So, the consumer is willing to give up 5.25 pounds of fish to obtain one more pound of steak during dinner.

19. U = 86x1 + 14x2 = 6159

When x1= 44,
86(44) + 14x2 = 6159
So, 3784 + 14x2 = 6159
So, 14x2 = 6159 - 3784 = 2375
So, x2 = 2375/14 = 169.64

Thus, x2 = 169.64